two angles are supplementary. if the larger is twice the smaller, what are the measures of the two angles? pls. use elimination
Answer by ewatrrr(24785)

 
Hi,
Let x and y represent the smaller and larger angle respectively
Question states***two angles are supplementary, larger is twice the smaller
x + y = 180° and y = 2x 0r 2x-y = 0
2x - y = 0
3x = 180°
x = 60° and y = 120°
Wish You the Best in your Studies.